OpenStudy (anonymous):

A block is placed at the bottom of the incline and given an initial velocity Vo directed along the incline. The coefficient of the kinetic friction between the block and the surface is 'm'. Find what should be the angle of the incline 'a' so the block goes the minimal distance along the incline. Find that distance

OpenStudy (anonymous):

 ok firt symbol for mas we use capital M and for acceleration we use t ok there is two forces so net force is mMg-Mgsina=Mt g(m-sina)=t now we kne v^2=u^2 + 2as 0=vo^2 +2ts vo^2/2g(m-sina)=s now for angle we can also find by g(m-sina)=t sina=-(t-gm)/g a=sin-(gm-t)/g